# Compactifications of phylogenetic systems and electrical networks

### v3

@@ 4.2. Compactifying Split Systems
Trees are a special case of both circular electrical networks and split systems. Kim [31] introduced a compactification by allowing edges of trees to become infinite in weight, resulting in a space of “phylogenetic oranges” [36]. Since resistance of $\infty$ corresponds to conductance of 0, the compactification of split systems will correspond to cactus networks.
@@
An analogous procedure can be followed for weighted split systems. Consider the dual polygon representation of a system in Figure 10(a). When we compactify this network by allowing an edge (say the diagonal ‘$e$’ labeled in green) to become infinite in weight, this corresponds to contracting the edge to a vertex, where the ends of any edges intersecting $e$ are also contracted along with $e$; see part (b). Note that this is dual to identification in cactus networks since resistance is the reciprocal of conductance. If this contraction results in overlapping of edges, the weights of these edges are summed. This is observed for edges $g+b+d$ in Figure 10(a) as well as $a+b$ and $d+f$ in part (b).
@@
The space $\mathbf{\Psi}_{n}$ is the set of weighted cactus circular split systems with clockwise cyclic order of $[n]$. Thus it is the set of Kalmanson metrics making up the blocks of $n\times n$ dissimilarity matrices $W$, where blocks correspond to the parts of the non-crossing partition of $[n]$.
The unweighted cactus split systems on $[n]$ correspond to the cells of a CW-complex structure on $\mathbf{\Psi}_{n}$. Containment of cells corresponds to either deletion or contraction (via infinity) of splits. The dimension of the top-dimensional cells of $\mathbf{\Psi}_{n}$ is (${n\choose 2}$, since that is the maximum number of splits compatible with a single cyclic order.
@@
Figure 9(b) shows the CW-complex structure of the space $\mathbf{\Psi}_{3}$, whereas Figure 14 shows the 1-skeleton of $\mathbf{\Psi}_{4}$: vertices in blue, edges in red. The $f$-vector for $\mathbf{\Psi}_{4}$ is (14, 28, 29, 24, 15, 6, 1) with 117 total cells. The 0-cells are counted by the Catalan numbers, since they correspond to the non-crossing partitions. Formulas for the total numbers of cells are given by Theorem 17 and its corollaries.
@@ 4.3. Maps
We extend the maps defined in Section 3 to take cactus networks as input. The process is largely straightforward, but we need to carefully define the Kron reduction and response/resistance matrices of a cactus network in order to see the operation of $\mathbf{\xi}$ and $\mathbf{\rho}$.
@@
In [33] is is stated that “There is a natural notion of the response matrix of a cactus network: we specify voltages at boundary vertices such that identified vertices are assigned the same voltage.” This matrix $M(N)$ can be posited as the limit of the response matrix as some of the edges of a network are taken to zero resistance. Also the Kron reduction of $N$ can be seen as expanded in a single disk to show edges with infinite conductance. Note that when there is an edge of $N$ incident on a (compactified) vertex (with multiple labels), the edges in $K(N)$ — and thus the entries in $M(N)$ — must evenly divide that conductance. Moreover, the response matrix of a cactus network will have entries of $\infty$ for the conductance between nodes that share a vertex, and corresponding diagonal entries of .$\infty$.- ∞ . The corresponding resistance matrix will have entries of 0 for pairs of nodes that share a vertex of the cactus network, but the other entries can be calculated directly by isolating the two nodes and using Kirchoff and Ohm’s laws to find effective resistance between them.
Example.
Figure 11 shows a cactus network $N$ and its dual $N^{*}$ together with their associated split systems. Each bulb of $N$ is its own Kron reduction. Also we can find a limiting $M(N)$ and $W(N)$ for the entire network, while here $M(N^{*})$ and $W(N^{*})$ exhibit block matrix structure.
@@
$M(N^{*})=\begin{bmatrix}-1&1&0&0&0&0\\ 1&-1&0&0&0&0\\ 0&0&-2/3&1/3&1/3&0\\ 0&0&1/3&-2/3&1/3&0\\ 0&0&1/3&1/3&-2/3&0\\ 0&0&0&0&0&0\\ \end{bmatrix}\hskip 50.58878ptW(N^{*})=\begin{bmatrix}0&1&\infty&\infty&\infty&\infty\\ 1&0&\infty&\infty&\infty&\infty\\ \infty&\infty&0&2&2&\infty\\ \infty&\infty&2&0&2&\infty\\ \infty&\infty&2&2&0&\infty\\ \infty&\infty&\infty&\infty&\infty&0\\ \end{bmatrix}$
The extension of the induced map $\mathbf{\sigma}$ gives us the ability to define it for inputs $N$ of both cactus networks and cactus networks with multiple connected components: We simply treat each connected component of $N$ separately. The connected components of $N$ give us a partition of $[n]$ with parts the boundary nodes of each component; this is the partition for $\mathbf{\sigma}(N)$. The splits of $\mathbf{\sigma}(N)$ on each part are those displayed in that component, with weights calculated via the spanning trees and 2-groves of that component using the edge weights (conductances). [^9] The multiple labels on nodes make no difference to the weight of a grove. We get the following:
Theorem 16.
For cactus network $N$, the map $\mathbf{\xi}$ coincides with both the Kalmanson and the induced split systems of the planar dual. That is, $\mathbf{\xi}(N)=\mathbf{\rho}(N^{*})=\mathbf{\sigma}(N^{*}).$ Respectively, we have $\mathbf{\rho}(N)=\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*}).$
Proof.
A cactus network $N$ has bulbs separated by vertices which are the boundary nodes sharing several labels from $[n]$. No dual edge can connect nodes from separate bulbs. Thus the dual $N^{*}$ has connected components that correspond to the bulbs of $N$. Vice-versa, connected components of $N$ correspond to bulbs in the dual $N^{*}$.
First then, the mapping of combinatorial structure can be treated just as for the single-bulb (ordinary) circular planar electrical networks, one bulb of the cactus network at a time. The same is true for electrical networks with more than one connected component: each component can be seen to contribute its own collection of splits on its own boundary nodes. For $N$ (or $N^{*}$) a cactus network, we can construct small response matrices separately for each bulb of $N$, choosing a name for the identified nodes on the boundary of that bulb. Then each of those bulbs will also have a resistance matrix of the same size. The maps $\mathbf{\xi}^{\prime}(N^{*})$ and $\mathbf{\rho}(N^{*})$ can thus operate bulb-wise. To see that this operation gives the expected values of the weights of the splits, we note that the process of compactifying a split system (via treating some splits as infinite) sums the weights of the splits which coincide after collapsing any edges to a point; see Figure 10. Thus the weight of the final split in the cactus split system $\mathbf{\xi}(N)$ is the same as the edge in the Kron reduction of a single bulb, which is the weight of the split found by $\mathbf{\rho}(N^{*})$ as in the connected case (Theorem 13). ∎
Figure 13. The 1-skeleton of $\mathbf{\Omega}_{4}$. Each 0-cell has 3,4, or 6 attached labeled 1-cells.
5. Enumeration
@@ 5.1. Lagrange Inversion
Our graphical map $\mathbf{\xi}$ takes any cactus circular network to a cactus planar split system. We leverage this map and show how to count these systems, all based on the following result:
@@
$A(x)\ =\ \left(\frac{1}{x}\right){\textup{Inverse}}\left(\frac{x}{B(x)}\right),\ \ \ \textup{where}\ \ \ B(x)\ =\ \sum_{k=0}^{\infty}2^{k\choose 2}x^{k},$
@@
To find the number of cactus split systems on $n,$ use the Lagrange inversion theorem on $x/B(x)$: Find the series expansion for $(B_{n}(x))^{n+1}$ (where $B_{n}(x)$ is the partial series up to $x^{n}$), take the coefficient of $x^{n}$ in that expansion, and divide by $n+1$. For instance, when $n=4$, we see
@@
where $585/5=117$, the coefficient of $x^{4}$ (585) divided by $n+1$ (5) results in the fourth term (117) of sequence A136654.
@@
We start by associating a species $A$ to the combinatorial structures of cactus split systems. Define $A([n])$ to be the set of pairs made up of a cactus split system on $[n]$ and a permutation on $[n].$ The permutation can be thought of as a 1-1 mapping of the exterior labels in their cyclic order to a new set of numerical tags $\{1,\dots,n\}$. This allows us to consider $A(x)$ as simultaneously the ordinary generating function of the numbers of structure (types) of cactus split systems and the exponential generating function of the species $A.$
Next, any instance of this species on $[n]$ can be created with the following steps. First we choose $k$ tags: any subset of $[n]$ of size $k$, and also a subset $S_{1}$ of our original cyclic labels of size $k$ but required to contain the label 1.101010For instance, to make the structure in Figure 12 we let $k=2$ and chose $S_{1}=\{1,15\}.$ The tags are not shown in Figure 12 because we only wanted to display the structure type. The tags are attached to the labels in $S_{1}$ (in any order) and a set of splits is assigned to $S_{1}.$ There are $2^{k\choose 2}$ possibilities for those splits.
Set $S_{1}$ is a part of a noncrossing partition of the circular labels $[n]$, and the remaining labels not in $S_{1}$ are subdivided by the contiguous portions of $S_{1}$ into at most $k$ contiguous (in the cyclic order) sets. These sets are then used to continue the process: the species $A$ is applied to each of them recursively, with all the possible partitions of the remaining $n-k$ tags. Thus we see a functional equation for the species:
@@
or equivalently, $A=B\circ(X\cdot A)$, where $B$ is the species of ordinary split systems and $X$ is the singleton species. Then in terms of exponential generating functions, $A(x)=B(xA(x))$ Multiplying both sides by $x$ and rearranging yields
$x\ =\ \frac{xA(x)}{B(xA(x))}\,,$
@@
$xA(x)\ =\ {\textup{Inverse}}\left(\frac{x}{B(x)}\right)$
@@ 6. Global Compactified Spaces
With future work in mind, we need to extend the definition of the space of circular electrical networks and their compactifications to encompass all circular permutations of $[n]$. We call this extending to the global case. Here we just introduce the spaces and enumerate some cells for small dimensions. Maps of this paper extend to the global spaces : but only $\mathbf{\sigma}$ and $\rho$ are well defined on the new global equivalence classes. The map $\mathbf{\xi}$ requires all its inputs and output to have a fixed cyclic ordering of nodes, so it is only well defined within a single chamber of the global space.
@@
Let $\mathbf{\Psi}^{global}_{n}$ denote the space of Kalmanson metrics on $[n]$, which bijectively correspond to circular split systems.
In [11] these systems are studied with the additional requirement that all trivial splits are present, which reflects a phylogenetic situation in which all taxa are temporally removed from common ancestors. That source points out that visually the diagrams of the split system can be twisted along splits that are either not crossed by another split or not present in the system—and still represent the same split system. When we compactify the space of split systems with a given cyclic order we represented the new points as non-crossing partitions of $[n]$, with the splits on each part of the partition. Now however we allow any partition of $[n]$, and any cyclic order for the polygon on each part of the partition.
@@
Let $\mathbf{\Psi}^{global}_{n}$ denote the space of Kalmanson metrics on parts of partitions of $[n]$, which correspond bijectively to circular split systems on each part of each partition.
Unweighted cactus networks with any cyclic order correspond to the cells of the resulting CW-complex structure on $\mathbf{\Omega}^{global}_{n}$. The 1-skeletons of $\mathbf{\Omega}^{global}_{n}$ and $\mathbf{\Psi}^{global}_{n}$ are made of 0-cells that have no edge (respectively no split) and 1-cells that have a single edge (single split). Thus these are identical as graphs. Figure 16 shows the 1-skeleton of $\mathbf{\Omega}^{global}_{4}$. Figure 17 shows the 1-skeleton of $\mathbf{\Psi}^{global}_{4}$, and is drawn as the mirror image of Figure 16: the map $\mathbf{\sigma}$ takes electrical networks in Figure 16 to the corresponding split system in the mirror image. Of course the number of higher dimensional cells in $\mathbf{\Omega}^{global}_{n}$ is generally fewer than in $\mathbf{\Psi}^{global}_{n}$ since planarity of the circuit is required, as opposed to just planarity of the splits. The $f$-vector for $\mathbf{\Omega}^{global}_{4}$ is (15, 31, 33, 26, 16, 9, 3) with 133 total cells. The $f$-vector for $\mathbf{\Omega}^{global}_{5}$ is (52, 160, 270, 345, 375, 378, 340, 255, 150, 60, 12) with 2397 total cells. The general numbers of cells for these global spaces, counted here by hand, is an open question, and the first step in an exciting new study of the topology of the spaces themselves. In Table 3 we list some statistics of the examples of spaces in this paper, with OEIS entry names for the sequences that have proven formulas for all $n$. Certain sequences begin with the same terms, since the spaces coincide (and have only one chamber) for $n\leq 3$.
The sequence of numbers of 0-cells for $\mathbf{\Omega}^{c}_{n}$ is shown to be the Catalan numbers in [18]. The sequence of numbers of 0-cells for $\mathbf{\Psi}_{n}$ is clearly the Bell numbers, since those give the numbers of partitions of $n$. That in turn implies that the Bell numbers also count the numbers of 0-cells in $\mathbf{\Omega}_{n}^{global}.$
Figure 16. The 1-skeleton of $\mathbf{\Omega}^{global}_{4}$.
Figure 17. The 1-skeleton of $\mathbf{\Psi}^{global}_{4}$.
Space number of 0-cells total cells $f$-vector
$\mathbf{\Omega}_{2}=\mathbf{\Omega}_{2}^{global}=\mathbf{\Psi}_{2}=\mathbf{\Psi}^{global}_{2}$ 1 2 (1, 1)
$\mathbf{\Omega}_{2}=\mathbf{\Omega}_{2}^{global}=\mathbf{\Psi}_{2}=\mathbf{\Psi}_{2}^{global}$ 2 3 (2, 1)
$\mathbf{\Omega}_{3}=\mathbf{\Omega}_{3}^{global}=\mathbf{\Psi}_{3}=\mathbf{\Psi}^{global}_{3}$ 1 8 (1, 3, 3, 1)
$\mathbf{\Omega}_{3}=\mathbf{\Omega}_{3}^{global}=\mathbf{\Psi}_{3}=\mathbf{\Psi}^{global}_{3}$ 5 15 (5, 6, 3, 1)
$\mathbf{\Omega}_{4}$ 1 52 [A111088] (1, 6, 14, 16, 10, 4, 1)
$\mathbf{\xi}(\mathbf{\Omega}_{4})$ 1 49 [Cor. 22] (1, 6, 14, 16, 9, 2, 1)
$\mathbf{\Psi}_{4}$ 1 64 [A006125] (1, 6, 15, 20, 15, 6, 1)
$\mathbf{\Omega}_{4}$ 14 [A000108] 105 [A001147] (14, 28, 28, 20, 10, 4, 1)
$\mathbf{\xi}(\mathbf{\Omega}_{4})=\mathbf{\sigma}(\mathbf{\Omega}_{4})$ 14 [A000108] 102 [A136654] (14, 28, 28, 20, 9, 2, 1)
$\mathbf{\Psi}_{4}$ 14 [A000108] 117 [A136654] (14, 28, 29, 24, 15, 6, 1)
$\mathbf{\Omega}^{global}_{4}$ 1 70 [open] (1, 6, 15, 20, 16, 9, 3)
$\mathbf{\Psi}_{4}^{global}$ 1 112 [open] (1, 7, 21, 34, 31, 15, 3)
$\mathbf{\Omega}^{global}_{4}$ 15 [A000110] 133 [open] (15, 31, 33, 26, 16, 9, 3)
$\mathbf{\sigma}(\mathbf{\Omega}^{global}_{4})$ 15 [A000110] 124 [open] (15, 31, 33, 26, 13, 3, 3)
$\mathbf{\Psi}_{4}^{global}$ 15 [A000110] 169 [open] (15, 31, 36, 38, 31, 15, 3)

### v4

@@ 4.2. Compactifying Split Systems
Trees are a special case of both circular electrical networks and split systems. Kim [32] introduced a compactification by allowing edges of trees to become infinite in weight, resulting in a space of “phylogenetic oranges” [37].
@@
An analogous procedure can be followed for weighted split systems. Consider the dual polygon representation of a system in Figure 10(a). When we compactify this network by allowing an edge (say the diagonal ‘$e$’ labeled in green) to become infinite in weight, this corresponds to contracting the edge to a vertex, where the ends of any edges intersecting $e$ are also contracted along with $e$; see part (b). Note that this is dual to identification in cactus networks since resistance is the reciprocal of conductance. If this contraction results in overlapping of edges, the weights of these edges are summed. This is observed for edges labeled $g+b+d$ in Figure 10(b) as well as $a+b$ and $d+f$.
@@
The space $\mathbf{\Psi}_{n}$ is the set of weighted circular cactus split systems with clockwise cyclic order of $[n]$. Thus it is the set of Kalmanson metrics making up the blocks of $n\times n$ dissimilarity matrices $W$, where blocks correspond to the parts of the non-crossing partition of $[n]$.
The unweighted cactus split systems on $[n]$ correspond to the cells of a CW-complex structure on $\mathbf{\Psi}_{n}$. Containment of cells corresponds to either deletion (weight = 0) or contraction (weight = $\infty$) of splits. The dimension of the top-dimensional cells of $\mathbf{\Psi}_{n}$ is (${n\choose 2}$, since that is the maximum number of splits compatible with a single cyclic order.
@@
Figure 9(b) shows the CW-complex structure of the space $\mathbf{\Psi}_{3}$, whereas Figure 12 shows the 1-skeleton of $\mathbf{\Psi}_{4}$. The $f$-vector for $\mathbf{\Psi}_{4}$ is (14, 28, 29, 24, 15, 6, 1) with 117 total cells. The 0-cells are counted by the Catalan numbers, since they correspond to the non-crossing partitions. Formulas for the total numbers of cells are given by Theorem 17 and its corollaries.
Figure 11. The 1-skeleton of $\mathbf{\Omega}_{4}$. The 0-cells have dark blue borders.
@@ 4.3. Maps
Since resistance of $\infty$ corresponds to conductance of 0, the compactification of split systems will correspond naturally to cactus electrical networks. We extend the maps defined in Section 3 to take cactus networks as input. The process is largely straightforward, but we need to carefully define the Kron reduction and response/resistance matrices of a cactus network in order to see the operation of $\mathbf{\xi}$ and $\mathbf{\rho}$.
@@
Example.
Figure 13 shows a cactus network $N$ and its dual $N^{*}$ together with their associated split systems. Each bulb of $N$ is its own Kron reduction. Indexed by the nodes in each bulb, but allowing the matrix entries for identified nodes to coincide, we have:
$M(N)_{1,2,5,6}=\begin{bmatrix}-1&1\\ 1&-1\\ \end{bmatrix}\text{~{}~{}}\text{~{}~{}~{}and~{}~{}~{}}\text{~{}~{}~{}}M(N)_{2,3,4,5,6}=\begin{bmatrix}-2&1&1\\ 1&-2&1\\ 1&1&-2\\ \end{bmatrix}.$
4.3.1. Graphical
The graphical map $\mathbf{\xi}$ operates on a cactus network $N$ by first finding the bulb-wise Kron reduction $K(N)$. Then each terminal node label is shifted counterclockwise to become attached instead to the nearest arc of the boundary in that direction. There is a unique nearest arc for each label since we have included empty bulbs. Then the edges of $K(N)$ are reinterpreted as weighted splits of the new cactus split system $\mathbf{\xi}(N).$ There is also a clockwise version, the cographical map $\mathbf{\xi}^{\prime}.$
4.3.2. Induced
The induced map $\mathbf{\sigma}$ operates on a cactus network by ignoring the bulb structure and instead acting component-wise. Each connected component displays splits of its set of terminal nodes, counting all the identified node labels as separate elements of that set. Thus each connected component of $N$ will become a bulb of the cactus split system $\mathbf{\sigma}(N).$ That is, the connected components of $N$ give us a non-crossing partition of $[n]$ with parts the boundary nodes of each component; this is the partition for $\mathbf{\sigma}(N)$. The splits of $\mathbf{\sigma}(N)$ on each part are those displayed in that component, with weights calculated via the spanning trees and 2-groves of that component using the edge weights (conductances). The multiple labels on nodes make no difference to the weight of a grove.
4.3.3. Kalmanson
The Kalmanson map $\mathbf{\rho}$ also operates component-wise. Each connected component of $N$ has its own Kalmanson resistance matrix, where identified nodes are assigned 0 resistance between them. The associated circular split system will be a bulb of the cactus split system $\mathbf{\rho}(N).$ We get the following:
Theorem 16.
For a cactus network $N$, the map $\mathbf{\xi}$ coincides with both the Kalmanson and the induced split systems of the planar dual. That is, $\mathbf{\xi}(N)=\mathbf{\rho}(N^{*})=\mathbf{\sigma}(N^{*}).$ Respectively, we have $\mathbf{\rho}(N)=\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*}).$
Proof.
We prove the first triple inequality, and the other follows from the same treatment with appropriate clockwise and counterclockwise rotations. A cactus network $N$ has bulbs separated by vertices which are the boundary nodes sharing several labels from $[n]$. No dual edge can connect nodes from separate bulbs. Thus the dual $N^{*}$ has connected components each consisting of nodes from the same bulb of $N$. Vice-versa, connected components of $N$ correspond to bulbs in the dual $N^{*}$. Thus we ask if performing the operation of $\mathbf{\sigma}$ (or $\mathbf{\rho}$) on a connected component of the dual $N^{*}$ has the same combinatorial result as performing the operation of $\mathbf{\xi}$ on the corresponding bulb of $N.$ Indeed, the same arguments hold as for the case of a single bulb with a single connected component in Theorem 13: edges in each of the bulb-wise Kron reductions $K(N)$ correspond to splits displayed by $N^{*}$.
We also need to check that the respective operations give matching values for the weights of the splits. This also follows from Theorem 13, applied to the single bulb of $N$ and the corresponding connected component of $N^{*}$. The weight of the split in the cactus split system $\mathbf{\xi}(N)$ is the same as the edge conductance in the Kron reduction of a single bulb, which in turn is the weight of the split found by $\mathbf{\rho}(N^{*})$ and $\mathbf{\sigma}(N^{*})$ as in the non-cactus case.
4.4. Summary of Maps
This construction raises new questions. First, although here we restrict our study to planar networks with a given clockwise order of nodes, the map $\mathbf{\xi}$ can operate on any network, as long as the nodes are given a clockwise order first. The results of that mapping for non-planar examples are interesting objects for future study.
Secondly the map $\mathbf{\xi}$ carries the structure of duality on electrical networks to a new duality on their image in the circular split systems. We note that the map $\mathbf{\xi}$ is reminiscent of the $T$-duality map on decorated permutations [36]. Moreover, $\mathbf{\xi}$ carries an equivalence relationship on circular split systems — under twisting as defined by Devadoss and Petti [12] — to a new combinatorial equivalence on response matrices which rewires the (possibly non-planar) electrical networks. What is the physical meaning of this equivalence? Indeed, $\mathbf{\xi}$ shows the poset and CW cell structures on circular split systems as a coarsening of those on circular planar networks. It also carries the cell structure of split systems to a new cell structure on all circular electrical networks. This new combinatorial structure is unexplored.
@@ 4.4. Summary of Maps
In [35] is is stated that “There is a natural notion of the response matrix of a cactus network: we specify voltages at boundary vertices such that identified vertices are assigned the same voltage.” This matrix $M(N)$ can be posited as the limit of the response matrix as some of the edges of a network are taken to zero resistance. Moreover, the response matrix of a cactus network will have entries of $\infty$ for the conductance between nodes that share a vertex, and corresponding diagonal entries of .$\infty$.- ∞ . The corresponding resistance matrix will have entries of 0 for pairs of nodes that share a vertex of the cactus network, but the other entries can be calculated directly by isolating the two nodes and using Kirchoff and Ohm’s laws to find effective resistance between them.
From Figure 13 we can find a limiting $M(N)$ and $W(N)$ for the entire network.
@@
5. Enumeration of Species
@@ 5.1. Lagrange Inversion
Our graphical map $\mathbf{\xi}$ takes any circular cactus network to a cactus planar split system. We leverage this map and show how to count these systems, all based on the following result:
@@
$\mathbf{\Psi}(x)\ =\ \left(\frac{1}{x}\right){\textup{Inverse}}\left(\frac{x}{\mathbf{\Psi}(x)}\right),\ \ \ \textup{where}\ \ \ \mathbf{\Psi}(x)\ =\ \sum_{k=0}^{\infty}2^{k\choose 2}x^{k},$
@@
To find the number of cactus split systems on $n,$ use the Lagrange inversion theorem on $x/\mathbf{\Psi}(x)$: Find the series expansion for $(\mathbf{\Psi}_{n}(x))^{n+1}$ (where $\mathbf{\Psi}_{n}(x)$ is the partial series up to $x^{n}$), take the coefficient of $x^{n}$ in that expansion, and divide by $n+1$. For instance, when $n=4$, we see
@@
Thus dividing the coefficient of $x^{4}$ by 4+1 gives us $|\mathbf{\Psi}_{4}|=585/5=117$.
@@
We start by associating a species $\mathbf{\Psi}$ to the combinatorial structures of cactus split systems  [^8] A great reference for this section is [3]. Define $\mathbf{\Psi}([n])$ to be the set of pairs made up of a cactus split system on $[n]$ and a permutation on $[n].$ The permutation can be thought of as a 1-1 mapping of the exterior labels in their cyclic order to a new set of numerical tags $\{1,\dots,n\}$. This allows us to consider $\mathbf{\Psi}(x)$ as simultaneously the ordinary generating function of the numbers of structure (types) of cactus split systems and the exponential generating function of the species $\mathbf{\Psi}.$
Next, any instance of this species on $[n]$ can be created with the following steps. First for any $k\leq n$ we choose $k$ tags: any subset $P_{1}\subset[n]$ of size $k$; and simultaneously we choose a size-$k$ starting subset $S_{1}$ of our original cyclic labels which is required to contain the least label 1. For instance, to make the structure in Figure 14 we let $k=2$ and chose $S_{1}=\{1,15\}.$ The tags are attached to the labels in $S_{1}$ (in any order) and a set of splits is assigned to $S_{1}.$ There are $2^{k\choose 2}$ possibilities for those splits. (The tags are not shown in Figure 14 because we only wanted to display the structure type.)
Now, the set $S_{1}$ is a part of a noncrossing partition of the clockwise circular labels $[n]$, and the remaining labels not in $S_{1}$ are subdivided by the contiguous portions of $S_{1}$ into at most $k$ contiguous (in the cyclic order) sets. These contiguous sets are then used to continue the process: the species $\mathbf{\Psi}$ is applied to each of them recursively using the least label as starting set, with all the possible partitions of the remaining $n-k$ tags. Each time, we draw the result as attaching the new cactus at the point on the old where the cyclic order of labels was cut. Thus we see a functional equation for the species:
@@
or equivalently, $\mathbf{\Psi}=\mathbf{\Psi}\circ(X\cdot\mathbf{\Psi})$, where $\mathbf{\Psi}$ is the species of ordinary split systems and $X$ is the singleton species. That is, any instance of the species of cactus split systems can be described recursively as an ordinary split system of cactus split systems each with at least one element (which we tag), but possibly with more: the attached cactus at the point after that tagged label. Therefore, in terms of exponential generating functions, $\mathbf{\Psi}(x)=\mathbf{\Psi}(x\mathbf{\Psi}(x))$ Multiplying both sides by $x$ and rearranging yields
$x\ =\ \frac{x\mathbf{\Psi}(x)}{\mathbf{\Psi}(x\mathbf{\Psi}(x))}\,,$
@@
$x\mathbf{\Psi}(x)\ =\ {\textup{Inverse}}\left(\frac{x}{\mathbf{\Psi}(x)}\right)$
@@ 5.3. Plabic Tilings and Ranges
The number of unweighted faithful circular split systems, which label the cells in the range of $\mathbf{\xi},$ is given by:
@@ 6. Global Compactified Spaces
With future work in mind, we need to extend the definition of the space of circular electrical networks and their compactifications to encompass all circular permutations of $[n]$. We call this extending to the global case. Here we introduce the spaces, enumerate some cells for small dimensions, and prove species formulas. Maps of this paper extend to the global spaces : but only $\mathbf{\sigma}$ and $\rho$ are well defined on the new global equivalence classes. The map $\mathbf{\xi}$ requires all its inputs and output to have a fixed cyclic ordering of nodes, so it is only well defined on equivalence classes within a single chamber of the global space.
@@
The cells of $\mathbf{\Omega}^{global}_{n}$ are labeled with minimal circular planar networks just as for $\mathbf{\Omega}_{n}$. The 0 and 1-dimensional cells are the same for both the global and clockwise spaces. For $n=4$, the extra cells of dimension 2–6 are seen in Figure 16. These can be added to columns 3–5 in Figure 5 to find the $f$-vector $(1,6,15,20,16,9,3)$ for $\mathbf{\Omega}^{global}_{4}$. The $f$-vector for $\mathbf{\Omega}^{global}_{5}$ is $(1,10,45,120,215,288,310,255,150,60,12)$.
Figure 16. The global circular planar electrical networks of $\mathbf{\Omega}^{global}_{4}$ that cannot be realized in planar clockwise cyclic order.
@@
Let $\mathbf{\Psi}^{global}_{n}$ denote the space of Kalmanson metrics on $[n]$, which bijectively correspond to circular split systems. This space is comprised of $(n-1)!/2$ chambers, one for each cyclic order, each a copy of $\mathbf{\Psi}_{n}.$
In [12] these systems are studied with the additional requirement that all trivial splits are present, which reflects a phylogenetic situation in which all taxa are temporally removed from common ancestors. That source points out that visually the diagrams of the split system can be twisted along splits that are either not crossed by another split or not present in the system—and still represent the same split system. The (known terms in) the sequence of total numbers of cells of $\mathbf{\Psi}^{global}_{n}$ are 2, 8, 112, 6976, 1332224, $\dots$. These are based on the hand-counted values in [12] and [40], each multiplied by $2^{n}$ to allow any subset of the trivial splits to be included in the split system. The general term is an open question.
When we compactify the space of split systems with a given cyclic order we represented the new points as non-crossing partitions of $[n]$, with the splits on each part of the partition. Now however we allow any partition of $[n]$, and any cyclic order for the polygon on each part of the partition.
@@
Let $\mathbf{\Psi}^{global}_{n}$ denote the space of Kalmanson metrics on parts of partitions of $[n]$, which correspond bijectively to circular split systems on each part of each partition. This space is comprised of $(n-1)!/2$ chambers, one for each cyclic order, each a copy of $\mathbf{\Psi}_{n}.$
Space number of 0-cells total cells $f$-vector $\chi$
$\mathbf{\Omega}_{2}=\mathbf{\Omega}_{2}^{global}=\mathbf{\Psi}_{2}=\mathbf{\Psi}^{global}_{2}$ 1 2 (1, 1) 0
$\mathbf{\Omega}_{2}=\mathbf{\Omega}_{2}^{global}=\mathbf{\Psi}_{2}=\mathbf{\Psi}_{2}^{global}$ 2 3 (2, 1) 1
$\mathbf{\Omega}_{3}=\mathbf{\Omega}_{3}^{global}=\mathbf{\Psi}_{3}=\mathbf{\Psi}^{global}_{3}$ 1 8 (1, 3, 3, 1) 0
$\mathbf{\Omega}_{3}=\mathbf{\Omega}_{3}^{global}=\mathbf{\Psi}_{3}=\mathbf{\Psi}^{global}_{3}$ 5 15 (5, 6, 3, 1) 1
$\mathbf{\Omega}_{4}$ 1 52 [A111088] (1, 6, 14, 16, 10, 4, 1) 0
$\mathbf{\xi}(\mathbf{\Omega}_{4})$ 1 49 [Cor. 22] (1, 6, 14, 16, 9, 2, 1) 1
$\mathbf{\Psi}_{4}$ 1 64 [A006125] (1, 6, 15, 20, 15, 6, 1) 0
$\mathbf{\Omega}_{4}$ 14 [A000108] 105 [A001147] (14, 28, 28, 20, 10, 4, 1) 1
$\mathbf{\xi}(\mathbf{\Omega}_{4})=\mathbf{\sigma}(\mathbf{\Omega}_{4})$ 14 [A000108] 102 [A136654] (14, 28, 28, 20, 9, 2, 1) 2
$\mathbf{\Psi}_{4}$ 14 [A000108] 117 [A136654] (14, 28, 29, 24, 15, 6, 1) 1
$\mathbf{\Omega}^{global}_{4}$ 1 70 [open] (1, 6, 15, 20, 16, 9, 3) 0
$\mathbf{\Psi}_{4}^{global}$ 1 112 [open] (1, 7, 21, 34, 31, 15, 3) 0
$\mathbf{\Omega}^{global}_{4}$ 15 [A000110] 133 [open] (15, 31, 33, 26, 16, 9, 3) 1
$\mathbf{\sigma}(\mathbf{\Omega}^{global}_{4})$ 15 [A000110] 124 [open] (15, 31, 33, 26, 13, 3, 3) 4
$\mathbf{\Psi}_{4}^{global}$ 15 [A000110] 169 [open] (15, 31, 36, 38, 31, 15, 3) 1